A escalator is moving downwards with a uniform speed u. A man of mass m is running upwards on it at a uniform speed v. If the height of escalator is h, the work done by man in going up the escalator is

Grade:11

1 Answers

Ashish Kumar
10 years ago
The relative velocity of the man w.r.t the the escalator = v - u. If the escalator was at rest then the man would have taken time t1 = h/v. But as the eslator is moving t2 =h/v -u.
The extra time taken , t2 -t1= hu/(v- u)v. The displacement in this extra time if the man was moving on a stationary escalator = (t2 - t1 ) v. Therefore the actual total displacement = h+huv /(v - u)v = hv / (v- u).
So work done = mghv/v- u





-858_signature1.GIF